Privacy and Security Notice

FEL Upgrade Project Weekly Brief
October 1-5, 2001

We continued our characterization measurements on the wiggler for the IR upgrade, uncovering some systematic
errors in the measurement apparatus.  We are pleased with wiggler characteristics measured thus far.  See the
wiggler section for details.

Members of the FEL project management team spent time with key managers and group leaders in the
Accelerator Division in a retreat this week to examine our present and future business lines.  Much of our time
was spent analyzing our capability to complete the 10 kW and 1 kW upgrade projects , the organization needed
to minimize conflicts and maximize benefits of our related projects in the CEBAF Upgrade and the SNS
accelerator, and the opportunities afforded by the potential 100 kW upgrade project and other future light source

A memo was distributed to the review team and invited guests for the next semiannual review for the FEL
Upgrade project on Nov. 8-9, soliciting particular agenda topics to supplement the project status reports.

WBS 4 (Injector):
Leak checked the ceramic stack and found 2 leaks in one kovar to ceramic braze. Modifying jigging for leak
checking the replacement ceramic prior to installation. The shop received the cathode support tube for machining.
The port tubes for the gun chamber have returned from being electropolished.

Gun HVPS - The design of the tank is still in progress.  The 3 toroids for the 3 stacks have been received

WBS 6 (RF):
Zone 3 - Only minor progress as the Lead Technician for this zone was at the Accelerator Retreat this week.

Zone 4 - All of the RF Control Modules have been calibrated and installed in the zone.  New CAMAC Stepper
Cards have been ordered to replace the missing ones.  They will be here in November.

Quarter - Equipment and hardware is being collected in order to mount the first solenoid and 100 kW klystron.
A lifting plate is being designed for the klystron.  The vendor, NWL, was contacted about making the 10 H
smaller.  A "Hold" was placed on its manufacture while its design is being checked.

WBS 8 (Instrumentation):
Progress continues on the new 32 channel Pico-motor relay chassis, the isolation module and a new Phase
Adjustable 30Hz trigger box for the IR cameras.

Firewire cables have been run to User Lab 1 and User Lab 3 to support new software using this technology.   A
GPIB repeater has been added to User Lab 1 for control of DMM in that lab during the up coming run.

5 new IR cameras have been modified and tested, 1 had a "twitch". One of those has been installed at Can 2 for
beam diagnostics at that location.

The new Areotech translation stages are being modified to conform with the existing stepper motor interface

Data Bases for the components of the Demo machine and the Upgrade machine are being developed for tracking
removal and storage of components as well as their subsequent re-installation.

Evaluation has begun on the Varian cold cathode guage controller that will be used to protect the new SNS SRF
input couplers.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  Consulted with Magnet Enterprises International on the details of the contract while the material is being
    prepared for a shipment next week
Injector Dipoles (DU/DV)
o  The drawings are in final review.
Arc Dipoles (GY, GX, GQ)
o  Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ)
    with the comments and corrections.
o  The Drawings of the 180° Dipole (DY) are being prepared for signature.
o  All steel necessary for the dipoles was received.
Magnet Measurement
o  Work continued on the dipole and quadrupole measurement stand design
3 inch quad (QX)
o  The assembly was awarded to New England Technicoil, the core vendor.
o  Work continues on the first article core.
o  64 coils are ready to ship from Magnet Enterprises International to the assembly vendor.
Trim Quad (QT)
o  This magnet remains in a procurement cycle.
Sextupole (SF)
o  Work continued on the magnetic model.
o  We will purchase one of the extra slabs from Bethlehem Steel left over from the dipole steel order for the solid
    cores of these sextupoles.
Octupole (OT)
o  DULY Research is working on a proposal to respond to a Task Order on the Engineering Services Contract
    to convert the magnetic model into a full blown magnetic design.
Corrector Dipoles (DB, DJ)
o  Milhous Co. has made approximately 20% of the coils.
Beam Line and Vacuum
o  The features of the chamber in the for the injection-reinjection beam line chambers were settled this week.
o  The girders and pedestals necessary for the return leg and the hardware for the entire machine entered remain
    in a procurement cycle.
o  Design work continues on the regions between the cryomodules.
o  Work continued on the injector line stands.

WBS 10 (Wiggler):
After some detective work, Tommy Hiatt and Michael Necaise found that the motor controller for the magnet
measurement stand was causing a 2.5mV offset on the probe.  This leads to the 5 G offset in the probe noted in
last weeks report.  The end corrector strength used last week was found to be the correct one.  Hiatt and
Necaise also found that a strange variation in the sinusoidal component in the vertical field was due to a 0.4
degree twist in the rail supporting the probe.  Since the horizontal field is so uniform however, it does not lead to a
large alteration of the horizontal trajectory.  Michael is trying several methods to eliminate the offset.

The final vacuum chamber welding is on track to be complete next week.  Both chambers have had the side
chambers welded on and one is out for final machining before welding on the flanges.  We will wait until fit-up to
weld up the two flanges that mate to the dispersion section chamber.

WBS 11 (Optics):
Procurements continue to be made on optical cavity components. These, and other components have transitioned
to detailing.  We are working on other components such as the insertable mirror assembly and the active
stabilization systems.  Copies of the presentation made at the internal review were sent to C. Brau and J.
Albertine for review and comments.

The upgrade near-Brewster windows were inspected, cleaned, and one was installed in user lab 1. New IR
cameras made by Raytheon are being installed in several locations in the optical transport, as well as on the larger
aperture pick off provided by the upgraded near-Brewster window in lab 1.  These cameras will give us
information on the beam position. (Kudos to the I&C group for developing and turning around a connector so we
could use the cameras so quickly.)

Some work has also been done in preparation for the next user run:

The protected-silver coated moly mirror in the collimator that absorbs the most power was found discolored.
This is a failure mode associated with running at high average power for extended periods of time, as was done
for the JTO run.  In an attempt to see whether we can mitigate this, we've switched to a protected-silver on
silicon mirror.  Meanwhile, the two damaged mirrors are being sent out to be polished and recoated.  The
aforementioned work in the labs and optical transport was done.  The lamps in the drive laser were replaced.

I attended the Boulder Damage Symposium, the premier conference devoted to optical damage in laser
components.  Two highlights that positively impact the upgrades to higher power are:

Magnetorheological finishing greatly reduces subsurface damage introduced during fabrication, with concomitant
increases in damage threshold and large reductions in damage probability near threshold.

New measurements on absorption losses in commercial grade sapphire are lower by an order of magnitude than
originally modeled.

We prepared for the start of the Fall user run next week by checking all susbsystems and running the gun through
a heat/clean recesiation cycle.  At week’s end after a 2 hr heat clean cycle we were not happy with the resulting
performance of the electrodes during HV conditioning nor the resulting final QE, so we instituted a second heat
clean/HV conditioning cycle for the weekend.